# Winning Big with Small Models: Knowledge Distillation vs. Self-Training for Reducing Hallucination in QA Agents

### v1

@@ Abstract
The deployment of Large Language Models (LLMs) in customer support is constrained by hallucination—generating false information—and the high cost of proprietary models. To address these challenges, we propose a retrieval-augmented question-answering (QA) pipeline and explore how to balance human input and automation. Using a dataset of questions about a Samsung Smart TV user manual, we demonstrate that synthetic data generated by LLMs outperforms crowdsourced data in reducing hallucination in finetuned models. We also compare self-training (fine-tuning models on their own outputs) and knowledge distillation (fine-tuning on stronger models’ outputs, e.g., GPT-4o), and find that self-training achieves comparable hallucination reduction. We conjecture that this surprising finding can be attributed to increased exposure bias issues in the knowledge distillation case and support this conjecture with post hoc analysis. We also improve robustness to unanswerable questions and retrieval failures with contextualized “I don’t know” responses. These findings show that scalable, cost-efficient QA systems can be built using synthetic data and self-training with open-source models, reducing reliance on proprietary tools or costly human annotations.  [^1] This work was conducted while Ashley Lewis was interning at Mitsubishi Electric Research Laboratories.
@@
for Reducing Hallucination in QA Agents
@@ 1 Introduction
  • We show that finetuning a model on its own generated answers (e.g., training Llama-3 on Llama-generated data) results in comparable hallucination mitigation to training it on GPT-4o-generated answers, despite GPT-4o being a more generally capable model.
@@ 3.2 Training Data
Table 1: FactScore results for the test set. Pretrained base models: Llama-3 and GPT-4o. Finetuned Llama-3-B models on the Nandy et al. (2021) dataset: Uncleaned (no data cleaning performed), Manual cleaned (cleaning done by the first author), AutocleanedL and AutocleanedG (cleaning done by Llama-3-70B and GPT-4o, respectively). Finetuned Llama-3-B models on synthetic data: SynthGPT (trained on data generated by GPT-4o), SynthLlama (trained on data generated by Llama-3-8B), and SynthLlama+ (same as SynthLlama, with additional negative examples).
@@ 3.3 Baseline and Experimental Models
We finetune Llama-3 on different variations of training data to analyze the effects of data source, cleaning method, and exposure bias on hallucination rates. Specifically, we train models on the following datasets using the Zheng et al. (2024) finetuning framework and parameters:
@@ 7 Limitations
Despite these insights, our study has limitations. First, our test set size is relatively small, particularly for human evaluation, where only 50 challenge and 50 regular test items were labeled. We did not want to overwhelm our annotators with too large of a task and judged that this was the maximum we could require. This limits the statistical power of our findings, making it difficult to detect smaller but meaningful performance differences. Expanding the evaluation set and conducting a larger-scale human evaluation in future work could provide a clearer picture of the impact of different training strategies.
@@ C.1 Answer Generation Prompt
The Bixby guide can be found by pressing the mic button once. The first time, a ’using Bixby’ button will appear. Click that for setup.

### v2

# Winning Big with Small Models: Knowledge Distillation vs. Self-Training for Reducing Hallucination in Product QA Agents

@@ Abstract
The deployment of Large Language Models (LLMs) in customer support is constrained by hallucination—generating false information—and the high cost of proprietary models. To address these challenges, we propose a retrieval-augmented question-answering (QA) pipeline and explore how to balance human input and automation. Using a dataset of questions about a Samsung Smart TV user manual, we demonstrate that synthetic data generated by LLMs outperforms crowdsourced data in reducing hallucination in finetuned models. We also compare self-training (fine-tuning models on their own outputs) and knowledge distillation (fine-tuning on stronger models’ outputs, e.g., GPT-4o), and find that self-training achieves comparable hallucination reduction. We conjecture that this surprising finding can be attributed to increased exposure bias issues in the knowledge distillation case and support this conjecture with post hoc analysis. We also improve robustness to unanswerable questions and retrieval failures with contextualized “I don’t know” responses. These findings show that scalable, cost-efficient QA systems can be built using synthetic data and self-training with open-source models, reducing reliance on proprietary tools or costly human annotations.
 [^1] This work was conducted while Ashley Lewis was interning at Mitsubishi Electric Research Laboratories.
@@
for Reducing Hallucination in Product QA Agents
@@ 1 Introduction
  • We show that finetuning a model on its own generated answers (e.g., training Llama-3 on Llama-generated data) results in comparable hallucination mitigation to training it on GPT-4o-generated answers, despite GPT-4o being a generally more capable model.
@@ 3.2 Training Data
Table 1: FactScore results for the test set. Pretrained base models: Llama-3 and GPT-4o. Finetuned Llama-3-8B models on the Nandy et al. (2021) dataset: Uncleaned (no data cleaning performed), Manual cleaned (cleaning done by the first author), AutocleanedL and AutocleanedG (cleaning done by Llama-3-70B and GPT-4o, respectively). Finetuned Llama-3-B models on synthetic data: SynthGPT (trained on data generated by GPT-4o), SynthLlama (trained on data generated by Llama-3-8B), and SynthLlama+ (same as SynthLlama, with additional negative examples).
@@ 3.3 Baseline and Experimental Models
We finetune Llama-3 on different variations of training data to analyze the effects of data source, cleaning method, and exposure bias on hallucination rates. Specifically, we train models on the following datasets using supervised fine-tuning (SFT) with LoRA adapters, following the parameters and framework of Zheng et al. (2024). During inference, we use greedy decoding with default settings:
@@ 4 Results and Analysis
Importantly, while the cleaning strategies evaluated here do not independently outperform the pretrained baseline, their primary utility lies elsewhere: enabling the generation of higher-quality synthetic QA data. As described in Section 4.2, models finetuned on synthetic data derived from cleaned examples (e.g., SynthLlama, SynthGPT) significantly outperform both manually and automatically cleaned models. This suggests that cleaning should be viewed not as an end in itself, but as a preparatory step for creating effective training data in low-resource settings.
@@ 7 Limitations
Despite these insights, our study has limitations. First, our test set size is relatively small, particularly for human evaluation, where only 50 challenge and 50 regular test items were labeled. We did not want to overwhelm our annotators with too large of a task and judged that this was the maximum we could require. This limits the statistical power of our findings, making it difficult to detect smaller but meaningful performance differences. Expanding the evaluation set and conducting a larger-scale human evaluation in future work could provide a clearer picture of the impact of different training strategies. Our work focuses on low-resource, domain-specific QA, reflecting common real-world settings—particularly in customer support—where large annotated datasets are rarely available. To our knowledge, the SmartTV corpus we use is the only publicly available product-manual QA dataset of its kind with a permissible license.
@@
Thirdly, we limit our experiments by using only Llama-3-8B as our base model. Our primary goal was to isolate the impact of training strategies—namely, self-training versus knowledge distillation—rather than compare model families. To ensure a fair comparison, we held the base model architecture constant across experimental conditions. Llama-3-8B was selected as a strong, cost-effective, and widely adopted open-source model. This choice supports reproducibility and reflects standard practice in related work; several recent papers on hallucination mitigation (e.g., Zhang et al. (2024) and Lin et al. (2024)) also restrict their experiments to only Llama-based models. However, future work with other architectures would be important to ensure generality of our findings here.
@@ C.1 Answer Generation Prompt
The Bixby guide can be found by pressing the mic button once. The first time, a ‘using Bixby’ button will appear. Click that for setup.